# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

### v3

@@ Abstract
Long-term causal effects often must be estimated from short-term data due to limited follow-up in healthcare, economics, and online platforms. Markov Decision Processes (MDPs) provide a natural framework for capturing such long-term dynamics through sequences of states, actions, and rewards. Double Reinforcement Learning (DRL) enables efficient inference on policy values in MDPs, but nonparametric implementations require strong intertemporal overlap assumptions and often exhibit high variance and instability. We propose a semiparametric extension of DRL for efficient inference on linear functionals of the $Q$-function—such as policy values—in infinite-horizon, time-homogeneous MDPs. By imposing structural restrictions on the $Q$-function, our approach relaxes the strong overlap conditions required by nonparametric methods and improves statistical efficiency. Under model misspecification, our estimators target the functional of the best-approximating $Q$-function, with only second-order bias. We provide conditions for valid inference using sieve methods and data-driven model selection. A central challenge in DRL is the estimation of nuisance functions, such as density ratios, which often entail difficult minimax optimization. To address this, we introduce a novel plug-in estimator based on isotonic Bellman calibration, which combines fitted $Q$-iteration with an isotonic regression adjustment. The estimator is debiased without requiring estimation of additional nuisance functions and reduces high-dimensional overlap assumptions to a one-dimensional condition. Bellman calibration extends isotonic calibration—widely used in prediction and classification—to the MDP setting and may be of independent interest.
@@
Keywords— Automatic debiasing, infinite-horizon MDPs, time-homogeneous dynamics, long-term causal inference, semiparametric restrictions, policy evaluation, isotonic calibration
1 Introduction
1.1 Motivation: Long-term causal inference
Randomized experiments—such as A/B tests and controlled trials—are widely used in healthcare, technology, and other industries to assess the impact of interventions on outcomes like survival, customer retention, and revenue. In industry, such experiments are often short-term due to practical constraints and the need for rapid decision-making. Consequently, analysts typically evaluate interventions using proxy metrics such as user engagement or click-through rates. These short-term outcomes inform decisions intended to improve long-term objectives. However, short-term experiments only yield unbiased estimates of short-term effects. This limitation has spurred growing interest in methods for inferring the long-term causal effects of policies from short-term data.
A common strategy for inferring long-term effects from short-term data involves surrogate methods. These approaches aim to link short-term experimental results to long-term outcomes by identifying intermediate variables, or “surrogates,” that are measured during the experiment (Athey et al., 2019). For example, a streaming platform may use engagement metrics—such as viewing hours or click rate—as surrogates for outcomes like annual membership retention. Given observational data containing long-term outcomes, these methods estimate the long-term causal effect by leveraging the surrogate–outcome relationship. A key assumption of surrogate methods is that the surrogate fully mediates the treatment effect. However, this assumption is violated when treatments involve sustained exposure beyond the experiment’s duration. For instance, evaluating a personalized recommendation algorithm—deployed continuously and adapting to user interactions—cannot rely on a short-term surrogate to capture the cumulative long-term effect, limiting the applicability of surrogate-based approaches.
These limitations motivate the use of dynamic modeling frameworks, which account for how long-term outcomes evolve in response to sequences of states, actions, and rewards under sustained treatment. For example, on a streaming platform, annual user retention under a new recommendation algorithm depends on daily engagement, shifting content preferences, and monthly subscription renewals. Recognizing these dynamics, Tran et al. (2023) proposed a method for estimating long-term effects of sustained treatments from short-term experiments, assuming that short-term observations sufficiently capture the long-term trajectory, even if they do not fully mediate the effect (see also Example 2 of Bibaut et al. (2021)). The core idea is to model the experiment’s temporal dynamics as a time-homogeneous Markov Decision Process (MDP) (Puterman, 1990) (see Figure 1), imposing Markov independence and stationarity on the state-action-outcome process. By linking long-term causal inference with offline reinforcement learning (Kaelbling et al., 1996), the authors develop nonparametric, efficient estimators of long-term treatment effects using Double Reinforcement Learning (van Der Laan et al., 2018; Kallus and Uehara, 2020, 2022).
Figure 1: DAG for trajectory under Markov Decision Process. The outcome $Y_{1}$, state $S_{2}$, and the trajectory need not be observed in the experiment.
Double Reinforcement Learning (DRL) enables statistically efficient inference on the value of a policy in nonparametric MDPs from off-policy trajectories. However, DRL faces fundamental challenges. First, DRL require sufficient overlap between the initial and future state distributions to ensure that all states relevant to the long-term trajectory are visited with positive probability. In high-dimensional or unbounded state spaces, this assumption is often violated (Mehrabi and Wager, 2024). Limited intertemporal overlap increases estimator variance, degrades stability, and necessitates large sample sizes. This challenge is analogous to, but distinct from, the poor performance of inverse propensity weighted estimators in cross-sectional studies with limited treatment overlap (D’Amour et al., 2021), as intertemporal state overlap cannot be ensured by randomization, given that future states are causally determined by prior states and actions. Second, DRL requires the estimation of complex nuisance functions via minimax optimization, which is computationally intensive and can be unstable in finite samples.
@@ 1.2 Contributions of this work
We develop a semiparametric extension of Double Reinforcement Learning (DRL) for inference on linear functionals of the $Q$-function—such as policy values—in infinite-horizon, time-homogeneous Markov Decision Processes (MDPs). The $Q$-function, which generalizes the regression function from static settings, encodes the expected cumulative outcome given the current state and action. Our approach addresses two key challenges in DRL: sensitivity to limited intertemporal overlap and the difficulty of estimating nuisance functions, such as density ratios. By imposing semiparametric restrictions on the $Q$-function, we relax the overlap conditions required for identification and improve the efficiency of resulting estimators.
Our key contributions are as follows:
  1. 1.
    We propose automatic debiased estimators for linear functionals of the $Q$-function—such as policy values—in time-homogeneous Markov Decision Processes (MDPs), under a semiparametric model for the $Q$-function. These estimators are doubly robust with respect to the estimation error of the $Q$-function and a certain Riesz representer of the linear functional. We further show how these estimators can be adjusted to enable valid inference for the best-approximating $Q$-function under model misspecification.
  2. 2.
    We show that misspecification of the $Q$-function incurs only second-order bias for the target functional, enabling valid inference under mild misspecification. We then provide conditions for valid inference using sieve methods and data-driven model selection.
  3. 3.
    We propose a novel debiased plug-in estimator for nonparametric inference that addresses two key challenges in DRL: (i) the computational burden and instability of minimax nuisance estimation, and (ii) sensitivity to model misspecification. The core of our approach is isotonic Bellman calibration, a generalization of isotonic calibration to MDPs, which may be of independent interest. We show that calibrating the $Q$-function estimator alone suffices to debias the plug-in estimator and provide valid nonparametric inference, without requiring additional nuisance estimation, such as density ratios.
Semiparametric restrictions have long been used to address instability in causal inference, particularly in cross-sectional settings. For example, nonparametric estimators of the average treatment effect (ATE), such as the augmented inverse probability weighted (AIPW) estimator (Robins et al., 1994), often suffer from high variance under limited treatment overlap. To mitigate this, prior work has leveraged working model assumptions (e.g., treatment effect homogeneity or partially linear models (Crump et al., 2006; Li et al., 2019; Robinson, 1988)) and dimension reduction techniques (Benkeser et al., 2020; D’Amour and Franks, 2021). Related ideas have been applied in data fusion for hidden confounding (Kallus et al., 2018), improving efficiency in randomized trials (van der Laan et al., 2024d), and enabling data-driven model selection (van der Laan et al., 2023). We extend these ideas to dynamic settings, addressing both treatment and intertemporal overlap, and the more complex nuisance estimation challenges that arise in MDPs.
Our work contributes to the growing literature on inference for off-policy evaluation (Murphy, 2003; Liu et al., 2018; Tang et al., 2019; Shi et al., 2022; Wang et al., 2023a), particularly through debiased and doubly robust estimation techniques (Tang et al., 2019; Kallus and Uehara, 2020; Shi et al., 2021; Kallus and Uehara, 2022; Mehrabi and Wager, 2024), as well as to research on long-term causal inference using MDPs and reinforcement learning methods (Liao et al., 2021; Tran et al., 2023; Nam et al., 2024). Existing approaches primarily target nonparametric, parametric, or sieve-based models of the $Q$-function, and focus on a specific linear functional: the policy value. We extend this literature by developing debiased inference procedures for general linear functionals of the $Q$-function under semiparametric restrictions. The challenge of limited intertemporal overlap was recently addressed by Mehrabi and Wager (2024), who proposed modified DRL estimators for policy values that adaptively truncate density ratios to improve performance when the ratio is unbounded but has finite variance. In contrast, our approach both relaxes overlap requirements and improves efficiency by imposing semiparametric structure on the $Q$-function. As a result, our methods remain valid even in settings where the density ratio has infinite variance or does not exist, rendering nonparametric identification impossible.
@@
This paper is organized as follows. Section 2 introduces Markov decision processes, $Q$-functions, and the target estimand. Section 3 presents our semiparametric DRL estimators, the corresponding asymptotic theory, and extensions. Section 4 discusses nuisance estimation strategies and associated challenges. In Section 5, we propose our Bellman-calibrated plug-in estimator. Finally, Section 6 presents numerical experiments.
@@ 2.1 Data Structure and Markov Decision Model
We consider a randomized experiment or observational study where participants sequentially receive treatments (actions) based on a policy. At each time $t$, a participant occupies a state $S_{t}$, which informs the choice of action $A_{t}$. The state-action pair determines an intermediate outcome $Y_{t}$, interpreted as an immediate reward or cost, and influences the transition to the next state $S_{t+1}$. Although participants are typically observed over a short time horizon, our goal is to estimate the long-term causal effects of a target policy $\pi$, which may differ from the behavior policy generating the observed data.
Formally, we represent an individual’s trajectory by the sequence $(S_{0},A_{0},Y_{0},S_{1},A_{1},Y_{1},S_{2},A_{2},\dots)\sim\mathbb{P}_{0}$, where states $S_{t}\in\mathcal{S}\subseteq\mathbb{R}^{d}$, actions $A_{t}\in\mathcal{A}$, and outcomes $Y_{t}\in\mathcal{Y}\subseteq\mathbb{R}$. The observed data consist of $n$ i.i.d. samples of a single state transition $(S_{0},A_{0},Y_{0},S_{1})$ from a distribution $P_{0}$ in a nonparametric model $\mathcal{P}$, forming the dataset $\mathcal{D}_{n}:=\{(S_{0,i},A_{0,i},Y_{0,i},S_{1,i}):i\in[n]\}$. This simplification entails little loss of generality, as multiple transitions from a single observation can be decomposed into individual transitions. Although such transitions may be dependent, our theoretical results extend using central limit theorems for Markov chains (Bibaut et al., 2021). We let $\pi(a\mid s)$ denote the (stationary) policy of interest, where $\pi(a\mid S_{t})$ represents the conditional probability of selecting action $a\in\mathcal{A}$ at time $t$ given the current state $S_{t}\in\mathcal{S}$. We denote by $\mathbb{P}_{0}$ the distribution of complete long-term trajectories induced by $P_{0}$, and by $P_{0,A_{0},S_{0}}$ the marginal distribution of $(A_{0},S_{0})$. To simplify notation, we write $S_{0}$ for any summary $S_{P_{0}}$ of the true distribution $P_{0}$.
@@
We assume the distribution $P_{0}$ of short-term observations $(S_{0},A_{0},Y_{0},S_{1})$ fully determines the distribution $\mathbb{P}_{0}$ of the long-term trajectory. Formally, we posit that the state-action-outcome process follows a time-homogeneous Markov decision process (Puterman, 1990), with data sequentially generated according to the nonparametric structural equation model (NPSEM) (Pearl, 2012):
$A_{t}:=f_{A}(S_{t},U_{A_{t}});\quad Y_{t}:=f_{Y}(A_{t},S_{t},U_{Y_{t}});\quad S_{t+1}$ $:=f_{S}(Y_{t},A_{t},S_{t},U_{S_{t+1}}),$
where $f_{A}$, $f_{Y}$, and $f_{S}$ are unknown deterministic functions, and the latent variables $\{U_{A_{t}},U_{Y_{t}},U_{S_{t+1}}\}$ are unobserved, mutually independent, and stationary random variables. This model imposes the Markovian assumptions that $A_{t}$ depends on its history only through $S_{t}$, $Y_{t}$ depends only on $A_{t}$ and $S_{t}$, and $S_{t+1}$ depends only on $Y_{t}$, $A_{t}$, and $S_{t}$. It also assumes stationarity over time, meaning that the conditional distributions $(S_{t+1}\mid Y_{t},A_{t},S_{t})$ and $(Y_{t}\mid A_{t},S_{t})$ are time-invariant. These assumptions become more plausible with a richer state space. For example, the state can be augmented to include multiple past time points, such as $\widetilde{S}_{t}:=(S_{t},S_{t-1},\dots,S_{t-k})$, or constructed using fixed, finite-dimensional summaries of historical information (van der Laan and Malenica, 2018).
The NPSEM allows us to define counterfactual MDPs, as the following example illustrates.
@@
2.2 Inferential objective and challenges in identification
In this work, we aim to estimate and perform inference on linear functionals of the $Q$-function (Kaelbling et al., 1996), a fundamental quantity connecting the causal effects of long-term policies to the observed short-term data. Formally, given a discount factor $\gamma\in[0,1]$, the $Q$-function associated with a policy $\pi$ is defined as the mapping:

### v4

@@ Abstract
Double Reinforcement Learning (DRL) enables efficient inference for policy values in nonparametric Markov decision processes (MDPs), but existing methods face two major obstacles: (1) they require stringent intertemporal overlap conditions on state trajectories, and (2) they rely on estimating high-dimensional occupancy density ratios. Motivated by problems in long-term causal inference, we extend DRL to a semiparametric setting and develop doubly robust, automatic estimators for general linear functionals of the $Q$-function in infinite-horizon, time-homogeneous MDPs. By imposing structure on the $Q$-function, we relax the overlap conditions required by nonparametric methods and obtain efficiency gains. The second obstacle—density-ratio estimation—typically requires computationally expensive and unstable min–max optimization. To address both challenges, we introduce superefficient nonparametric estimators whose limiting variance falls below the generalized Cramér–Rao bound. These estimators treat the $Q$-function as a one-dimensional summary of the state–action process, reducing high-dimensional overlap requirements to a single-dimensional condition. The procedure is simple to implement: estimate and calibrate the $Q$-function using fitted $Q$-iteration, then plug the result into the target functional, thereby avoiding density-ratio estimation altogether.
@@
Keywords— Double reinforcement learning, semiparametric inference, policy evaluation, long-term causal inference, automatic debiasing, superefficiency, calibration
1 Introduction and motivation
Randomized experiments such as A/B tests and clinical trials are widely used to assess the effects of interventions. In industry, econometrics, and clinical applications, these experiments are often short-term, even though the decisions they inform involve long-term outcomes such as survival, customer retention, or revenue. This gap has motivated a central question in recent research: How can we estimate the long-term causal effect of an intervention when only short-term experiments are available? (athey2019surrogate; van2023estimating; tran2023inferring; bibaut2023long; nam2024predicting; imbens2025long)
A common strategy is to use surrogate outcomes—intermediate variables observed during the experiment that are assumed to mediate the long-term effect (prentice1989surrogate; frangakis2002principal; athey2019surrogate). These methods combine short-term experimental data with long-term observational data by treating the surrogate as a proxy for the eventual outcome. For example, a streaming platform might use one-month watch time in an A/B test as a surrogate for annual subscription retention recorded in historical data. A key limitation of surrogate-based approaches is that the short-term surrogate must fully mediate the long-term treatment effect (vanderweele2013surrogate; weintraub2015perils; stijven2025evaluation). This assumption fails in settings where treatment effects accumulate over time (tran2023inferring). For instance, annual retention under a new recommendation algorithm depends on evolving user behavior that cannot be captured by a single short-term surrogate.
Figure 1: DAG for Markov Decision Process. $Y_{1}$ and $S_{2}$ need not be observed in the experiment.
These limitations motivate dynamic models that capture how long-term outcomes evolve in response to sequences of states, actions, and rewards under sustained treatment. To this end, tran2023inferring proposed estimating long-term effects from short-term experiments using a time-homogeneous Markov Decision Process (MDP) (Figure 1). Unlike surrogate-based methods, the MDP approach does not require the long-term effect to be mediated by a single short-term variable; instead, it models the full state–action trajectory. Earlier work in this direction includes battocchi2021estimating, liao2021off, and Example 2 of bibaut2021sequential. These approaches link long-term causal inference to offline reinforcement learning (kaelbling1996reinforcement) and enable efficient nonparametric estimation via Double Reinforcement Learning (DRL) (van2018online; kallus2020double; kallus2022efficiently).
DRL enables statistically efficient inference on policy values in nonparametric MDPs using off-policy data, but it faces two major practical challenges. First, DRL requires sufficient intertemporal overlap in state trajectories, meaning that the distribution of states observed early in the trajectory has support over the states that determine long-term outcomes. When this overlap is weak, DRL estimators become unstable and require prohibitively large samples—a difficulty that is especially severe in high-dimensional or unbounded state spaces (mehrabi2024off). Unlike standard policy overlap, intertemporal overlap is more restrictive because future states cannot be randomized: even perfect treatment randomization does not guarantee support in the state distribution over time. A closely related issue is well known in offline RL, where short-horizon behavior data do not cover the long-term state distribution induced by the target policy, leading to extrapolation error and instability under temporal distribution shift (buckman2020importance; szepesvari2022algorithms). Although value estimation under limited overlap has been widely studied, valid statistical inference in such settings remains largely unresolved (fujimoto2019off; levine2020offline; jin2021pessimism).
Second, DRL requires estimating high-dimensional nuisance components such as occupancy density ratios, typically via min–max optimization over adversarial function classes. State-of-the-art methods include DICE (nachum2019dualdice; zhang2020gendice) and minimax weight estimation approaches (liu2018breaking; uehara2020minimax; dikkala2020minimax). In practice, these adversarial estimators are computationally demanding, tuning-sensitive, and often unstable in finite samples (jiang2020minimax; kallus2020double). A common workaround is to restrict the function class used in the inner maximization (e.g., to linear or RKHS models), but this induces misspecification bias (kallus2020double; kim2022lobsdice; olivas2025source).
@@ 1.1 Contributions of this work
Motivated by practical challenges, we extend the DRL framework to incorporate semiparametric model restrictions and introduce a new class of superefficient estimators that address the twin challenges of limited intertemporal overlap and high-dimensional nuisance estimation. Our three main contributions are as follows.
Semiparametric and automatic inference for DRL. We develop a semiparametric framework for estimation and inference in DRL under model restrictions on the $Q$-function in time-homogeneous MDPs. The $Q$-function is the dynamic analogue of the outcome regression, linking long-term cumulative rewards to short-term data through the Bellman equation (bellman1952theory). By imposing structure on the $Q$-function, our approach relaxes the stringent intertemporal overlap conditions required by fully nonparametric methods and yields improved statistical efficiency. Related gains have been shown in the special case of linear MDPs, where both the transition kernel and reward function are assumed linear in a known feature map (xie2023semiparametrically). In contrast, we show that similar benefits arise under substantially weaker assumptions: it is sufficient to restrict only the $Q$-function, without imposing a parametric model on the underlying MDP. The intuition is that any restriction on the $Q$-function implicitly constrains both the reward function and the transition dynamics through the Bellman equation. Beyond policy values, we consider inference for general $L^{2}$-continuous linear functionals of the $Q$-function, and show that debiasing requires estimating only one additional nuisance component—the corresponding Riesz representer, which generalizes the discounted state–action occupancy ratio. We construct automatic DRL estimators whose form does not depend on the target functional and which are doubly robust to misspecification of both the $Q$-function and its Riesz representer.
Superefficient nonparametric estimators under limited overlap. When model restrictions are misspecified, DRL estimators may introduce bias, whereas fully nonparametric methods remain valid but can become unstable, highly variable, or nonregular under limited intertemporal overlap (imbensOverlapEstimand2006; benkeser2020nonparametric; mehrabi2024off). To address this, we develop a new class of superefficient nonparametric estimators for continuous linear functionals of the $Q$-function. These estimators are asymptotically normal with variance strictly below the generalized Cramér–Rao bound, enabling inference in irregular settings where root-$n$ regular estimation is infeasible, such as when occupancy ratios have infinite variance (mehrabi2024off). They achieve this by using the $Q$-function as a one-dimensional summary of the state–action space, thereby reducing high-dimensional overlap requirements to a single-dimensional condition.
Calibration removes the need to estimate density ratios. Existing DRL estimators rely on adversarial min–max optimization to estimate the discounted occupancy ratio, a step that is computationally intensive, tuning-sensitive, and unstable. Our proposed superefficient estimator bypasses this step entirely by using a calibrated plug-in estimator. The method is simple: estimate the $Q$-function using fitted $Q$-iteration, calibrate it, and plug the result into the target functional. We show that calibration implicitly carries out the necessary bias correction, obviating the need to estimate density ratios or, more broadly, the Riesz representer. Unlike existing DRL estimators, our approach requires only fitted $Q$-iteration (munos2008finite), making nuisance estimation computationally efficient and stable, and supported by well-developed theory (fan2020theoretical; hu2025fast; munos2008finite). To implement this idea, we introduce fitted $Q$-calibration, a dynamic extension of isotonic calibration (zadrozny2002transforming; niculescu2005predicting).
@@
This paper is organized as follows. Section 2 introduces Markov decision processes, $Q$-functions, and the target estimand. Section 3 presents our semiparametric DRL estimators, the corresponding asymptotic theory, and extensions. In Section 4, we propose our superefficient estimator and fitted $Q$-calibration. Finally, Section 5 presents numerical experiments.
1.2 Related work
Off-policy evaluation. Our work builds on recent advances in inference for off-policy evaluation (murphy2003optimal; liu2018breaking; tang2019doubly; shi2022statistical; wang2023projected), particularly debiased and doubly robust estimation methods (tang2019doubly; kallus2020double; shi2021deeply; kallus2022efficiently; mehrabi2024off). Existing approaches typically model the $Q$-function using nonparametric or parametric methods and focus on a single linear functional—the policy value. Parametric inference in linear MDPs has been studied by xie2023semiparametrically and wu2024statistical. We extend this literature by developing debiased inference procedures for general linear functionals of the $Q$-function under semiparametric restrictions. We focus on time-homogeneous MDPs, which are especially relevant for long-term causal inference and reinforcement learning, and which avoid the curse of the infinite horizon that arises in non-homogeneous models (liu2018breaking; kallus2022efficiently).
Addressing intertemporal overlap. Early work on nonparametric inference in off-policy evaluation required strong intertemporal overlap, assuming the occupancy density ratio was uniformly bounded (van2018online; kallus2020double; kallus2022efficiently). Later work relaxed this to a finite-variance assumption via ratio truncation (mehrabi2024off), but fully nonparametric estimators can still require large samples when overlap is weak, and the ratio itself may have infinite variance or fail to exist, making root-$n$ regular inference impossible. In contrast, we address overlap through model restrictions, and our superefficient estimator enables valid nonparametric inference even when the density ratio is ill-behaved. In value estimation and policy optimization, limited overlap is often handled by regularizing value estimates in poorly supported regions (yu2020mopo; liu2020provably; jin2021pessimism), but these methods do not extend naturally to statistical inference. Our use of semiparametric structure achieves a similar form of extrapolation from well-supported to weakly supported regions, but is designed for valid inference rather than policy optimization.
Semiparametric restrictions and efficiency. Semiparametric restrictions have long been used to stabilize causal inference in cross-sectional settings. For example, while fully nonparametric estimators such as AIPW (robinsCausal) may suffer from high variance under limited treatment overlap, estimators based on partially linear models are far less sensitive to overlap violations (imbensOverlapEstimand2006; li2019overlapWeights; robinson1988root). Related ideas appear in data fusion, where model restrictions are used to address unobserved confounding (kallus2018removing), and in randomized trials, where such restrictions improve efficiency by leveraging external data (van2024adaptive).
Superefficient estimation. Superefficient ATE estimators have been developed to permit nonparametric inference under limited overlap (CTMLElasso; d2021deconfounding; moosavi2023costs; van2023adaptive), typically through data-driven confounder selection or dimension reduction (shortreed2017outcome; HALoutcomeAdapt; benkeser2020nonparametric). Our work is closest to benkeser2020nonparametric, who use the outcome regression function as a one-dimensional summary for adjustment. We extend this idea to dynamic settings, showing that the $Q$-function plays an analogous role in MDPs. We achieve debiasing via isotonic calibration rather than targeted minimum loss estimation, which simplifies the procedure and avoids additional nuisance estimation.
Calibration and isotonic regression. Calibration has a long history in forecasting and machine learning as a tool for improving the reliability of classification and prediction models (lichtenstein1977calibration; platt1999probabilistic; zadrozny2001obtaining). Isotonic calibration (zadrozny2002transforming; niculescu2005predicting) is a widely used instance based on isotonic regression (barlow1972isotonic). In causal inference, calibration has primarily been applied to improve nuisance estimators, such as CATEs and propensity scores (gutman2022propensity; van2023causal; ballinari2024improving; van2024stabilized; rabenseifner2025calibrationstrategiesrobustcausal). This literature is distinct from causal isotonic regression (westling2020causal), a shape-constrained method for inference on monotone dose–response curves.
In contrast, we use calibration not as a nuisance-refinement step but as a core ingredient for inference. Our approach is motivated by the calibrated DML framework of vanderlaan2024automatic, which employs calibration to achieve doubly robust inference for linear functionals of a regression function, yielding estimators that remain asymptotically normal even when one of the nuisance components is inconsistently estimated. While we extend this idea to the MDP setting, our objective differs. Rather than targeting double robustness, we use calibration to construct a super-efficient plug-in estimator that avoids estimation of the Riesz representer. The dynamic setting also introduces new challenges: the $Q$-function is defined as the solution to a linear inverse problem, requiring new calibration techniques and corresponding theory. To our knowledge, calibration of $Q$-functions in MDPs has not previously been studied, and our dynamic generalization of isotonic calibration is novel and of independent methodological interest.
@@ 2.1 Data Structure and Markov Decision Model
We consider a randomized experiment or observational study in which individuals sequentially receive treatments (actions) according to a policy. At each time $t\in\{0,1,2,\dots\}$, an individual occupies a state $S_{t}\in\mathcal{S}$, which informs the choice of action $A_{t}\in\mathcal{A}$. The resulting state–action pair generates an intermediate outcome $Y_{t}\in\mathbb{R}$, interpreted as an immediate reward or cost, and determines the next state $S_{t+1}$. The observed data consist of $n$ i.i.d. samples of a single transition, $\mathcal{D}_{n}:=\{(S_{0,i},A_{0,i},Y_{0,i},S_{1,i}):i\in[n]\},$ drawn from a distribution $P_{0}$ in a nonparametric model $\mathcal{P}$. Although only one transition is observed per individual, datasets with multiple transitions can be decomposed into such units, and any remaining dependence can be handled using Markov chain limit theory (bibaut2021sequential). The goal is to perform long-term causal inference for a counterfactual policy $\pi(a\mid s)$, assumed stationary and admitting a density with respect to a measure $\nu$ on $\mathcal{A}$. For notational convenience, we write $f_{0}$ for any population quantity $f_{P_{0}}$ defined under $P_{0}$.
@@
The infinite-horizon trajectory $(S_{0},A_{0},Y_{0},S_{1},A_{1},Y_{1},S_{2},A_{2},\dots)$ is generated by a time-homogeneous Markov decision process with joint law $\mathbb{P}_{0}$ (puterman1990markov). Formally, we represent the data-generating mechanism using the nonparametric structural equation model (NPSEM) (pearl2012causal):
$A_{t}$ $=f_{A}(S_{t},U_{A_{t}}),\qquad Y_{t}$ $=f_{Y}(A_{t},S_{t},U_{Y_{t}}),\qquad S_{t+1}$ $=f_{S}(Y_{t},A_{t},S_{t},U_{S_{t+1}}),$
where $f_{A}$, $f_{Y}$, and $f_{S}$ are unknown deterministic functions, and the latent variables $\{U_{A_{t}},U_{Y_{t}},U_{S_{t+1}}\}$ are mutually independent, unobserved, and stationary. This encoding implies the Markov property: $A_{t}$ depends on the past only through $S_{t}$, $Y_{t}$ depends only on $(A_{t},S_{t})$, and $S_{t+1}$ depends only on $(Y_{t},A_{t},S_{t})$. We also assume time-homogeneity, so the conditional laws of $Y_{t}$ and $S_{t+1}$ do not depend on $t$. Under these assumptions, the full trajectory law $\mathbb{P}_{0}$ is determined by the distribution $P_{0}$ of the first transition. These assumptions become more plausible when the state space is sufficiently rich—for example, by augmenting the state with a fixed number of past values, $\widetilde{S}_{t}:=(S_{t},S_{t-1},\dots,S_{t-k})$, or with low-dimensional summaries of historical information (van2018robust). The NPSEM allows us to define counterfactual MDPs, as illustrated in the next example.
@@
2.2 Inferential objective
Our goal is to perform inference on $L^{2}$-continuous linear functionals of the $Q$-function, which links long-term causal effects to short-term data. For a given discount factor $\gamma\in[0,1]$, the $Q$-function associated with a policy $\pi$ is defined as